OpenStudy (anonymous):

In the slope-intercept equation of a line, which part of the equation gives the y-coordinate of the point where the line crosses the y-axis? A. The constant term B. The coefficient of the x-term C. The coefficient of the y-term D. The x-term

hartnn (hartnn):

slope intercept form : y = mx+c line crossing y axis ---> Y intercept when x= 0

hartnn (hartnn):

plug in x=0 in y=mx+c what u get?

OpenStudy (anonymous):

y= C +1?

OpenStudy (anonymous):

wait no

hartnn (hartnn):

waiting

OpenStudy (anonymous):

ok alright.... y=m0+c right?

hartnn (hartnn):

yes m*0 = 0

hartnn (hartnn):

so y=c right ? and 'c' is called the constant :)

hartnn (hartnn):

*the constant term

OpenStudy (anonymous):

@hartnn you might have learned this differently than me but I thought it was y=mx+b

hartnn (hartnn):

yes, it doesn't make a difference :)

OpenStudy (anonymous):

so the answer would be A?

hartnn (hartnn):

yes :)

OpenStudy (anonymous):

oh ok I was just confused but that does help understand the problem

OpenStudy (anonymous):

Yay :)

OpenStudy (anonymous):

thanks for the help!

hartnn (hartnn):

welcome ^_^